# Neural network-augmented eddy viscosity closures for turbulent premixed jet flames

### v1

@@ 5.1 Computational cost of adjoint-based training and inference
5.2 Comparison to alternative optimization methods
Several key differences inform the choice of PDE-constrained, embedded optimization (DPM) over standard data-driven, a priori optimization methods. While DPM’s training costs are higher (due to the need to make predictions and solve adjoint equations), it offers several distinct advantages for a posteriori predictions, which ultimately are the objective of most modeling efforts.
First, DPM achieves model consistency [32] during training by accounting for the feedback that occurs between the trained model and the governing PDEs during predictions. In contrast, a priori optimization does not incorporate the feedback that occurs in a posteriori predictions, which can lead to poor predictive performance and even instability [31, 33].
@@ 5.2 Comparison to alternative optimization methods
Second, DPM permits objective functions to be obtained from any PDE solution variables, including derived quantities such as spatially filtered and/or time-averaged statistics [33]. This enables the DPM optimization procedure to target experimental data, from which the closure terms needed by a priori optimization may not be readily available, or even data from “higher-order” simulations such as direct simulation Monte Carlo solutions of the Boltzmann equation [36], which might lack obvious connections to the modeled closure terms. In contrast, a priori optimization must directly target the outputs of the closure model and hence is typically limited to DNS target data.
@@
The main challenge of PDE-constrained optimization (not necessarily limited to DPM) is the cost of solving the optimization problem by iterating over the forward PDEs. DPM uses adjoint equations to obtain the gradients needed for optimization, which significantly alleviates the compute and memory requirements of non-graph-breaking alternatives (e.g., [51]). The adjoint equations can be complicated, especially for high-dimensional problems and detailed chemical kinetics. However, modern algorithmic differentiation frameworks, such as that provided by PyTorch [43], have greatly simplified the process of constructing the discrete-exact adjoint PDEs and have made DPM accessible for large-scale flow simulations.

### v2

@@ Abstract
Novelty and Significance
We develop a deep learning turbulence closure method for RANS calculations of turbulent premixed flames. The closure method embeds an untrained neural network into the RANS equations and then optimizes it over the flow solution using an adjoint-based technique. Novelty: This is the first application of solver-embedded deep learning to turbulent premixed jet flames. Significance: The method is a new, general-purpose closure-modeling framework for turbulent flames. For the present turbulent premixed jet flames, the method significantly reduces the error of a posteriori RANS predictions. The trained models generalize this improved accuracy across a wide range of Damköhler numbers, even in combustion regimes that are far out-of-sample from those used to train a particular model, which is not typical of deep learning closures.
Author Contributions
  • 1.
    Priyesh Kakka: conceptualization, methodology, software, validation, formal analysis, investigation, data curation, writing - original draft, visualization
  • 2.
    Jonathan F. MacArt: conceptualization, software, resources, writing - review & editing, supervision, project administration, funding acquisition
@@ 5.1 Computational cost of adjoint-based training and inference
5.2 Comparison of DPM to a priori models
We assess the trade-offs between PDE-constrained embedded optimization (DPM) and standard a priori (offline) modeling in two steps. As a “best-case” scenario, we first evaluate the performance of RANS using $\mu_{t}$ obtained directly from the DNS-evaluated $k$ and $\epsilon$. Such an approach would clearly be infeasible for general-purpose RANS, as it requires DNS data for the configuration of interest. However, it permits evaluation of model-consistency errors (due to the fact that the a priori model is not optimized consistently with the RANS equations) independently of any model-form errors (for example, inadequacy of a neural network fit to data). Second, we optimize a neural network model to the a priori, DNS-evaluated $\mu_{t}$ as an example of a generalizable a priori modeling approach.
We first evaluate the turbulent viscosity directly using the a priori, DNS-evaluated TKE and dissipation rate using (30); we designate this closure $\mu_{t}^{k\epsilon}(Q_{\text{{dns}}};C_{\mu})$ with $C_{\mu}=0.09$ as before. The resulting eddy viscosity is comparable in magnitude to the DPM-optimized, ME0-modeled $\mu_{t}(Q;\theta)$, shown in Figure 17 for $\mathrm{Da}_{s}=0$. The a priori, DNS-evaluated turbulent thermal diffusivity $\mu_{t,E}^{k\epsilon}(Q_{\text{{dns}}};C_{\mu})$ is computed using $\mathrm{Pr}_{t}=0.9$.
Figure 17: Comparison of DPM-optimized eddy viscosity and eddy diffusivity using the ME0 model to a priori, DNS-evaluated $k$$\epsilon$ eddy viscosity. Shown for $\mathrm{Da}_{s}=0$ at $t=15$.
We additionally train a neural network for the a priori eddy viscosity using a formulation analogous to the DPM model (31),
$\mu_{t,ap}(\theta)\leftarrow\mathcal{N}_{ap}\left(\frac{\partial\bar{\rho}}{\partial y},\,\frac{\partial\bar{\rho}\widetilde{u}}{\partial y},\,\frac{\partial\bar{\rho}\widetilde{v}}{\partial y},\,\frac{\partial\bar{\rho}\widetilde{E}}{\partial y},\,\frac{\partial\bar{\rho}\widetilde{Y}_{P}}{\partial y},\,\frac{\partial\bar{p}}{\partial y},\,\mathrm{Da}_{s}\,\dot{\omega}_{P}(\overline{\rho},\widetilde{Y_{P}},\widetilde{T}),\,t\right),$ (39)
with the turbulent thermal diffusivity likewise obtained using $\mathrm{Pr}_{t}=0.9$. The architecture of $\mathcal{N}_{ap}$ is the same as that of the DPM networks. We train $\mathcal{N}_{ap}$ over 7,000 DNS time steps across 2,000 epochs, which requires approximately 3 hours on a single NVIDIA RTX A2000 GPU. The training employs the MSE loss
$J(\theta)=\frac{1}{N^{\prime}_{y}}\sum_{i=1}^{N^{\prime}_{y}}\left(\left(\mu_{t,ap}(Q_{\text{{dns}}};\theta)-\mu_{t}^{k\epsilon}(Q_{\text{{dns}}};C_{\mu})\right)^{2}\right)(y_{i}).$ (40)
The resulting $\mu_{t,ap}(Q;\theta)$ closes the RANS equations. As with any a priori model, inference in a posteriori simulations uses the predicted $Q$ rather than the DNS-evaluated $Q_{\text{{dns}}}$ used for training, which is a potential source of model-consistency error.
Figure 18 compares the errors in the a posteriori RANS temperature and cross-stream velocity, with respect to the DNS target data, using the $k$$\epsilon$ model, DPM-optimized ME0 model, direct DNS-evaluated $\mu_{t}^{k\epsilon}(Q_{\text{{dns}}};C_{\mu})$, and a priori, neural network-modeled $\mu_{t,ap}(Q;\theta)$ for the baseline $\mathrm{Da}_{s}=0$ case. For both a priori closures, the in-sample errors are significantly higher than those of the DPM-optimized ME0 model.
Table 3 lists the cumulative MSE errors for these four models for the in-sample $\mathrm{Da}_{s}=0$ case and out-of-sample predictions at $\mathrm{Da}_{s}=20{,}000$. In general, RANS predictions using the direct DNS-evaluated $\mu_{t}^{k\epsilon}$ are more accurate than the corresponding $k$$\epsilon$ RANS, as could be anticipated for the in-sample $\mathrm{Da}_{s}=0$ case, though this also holds when the $\mathrm{Da}_{s}=0$ DNS-evaluated eddy viscosity is used for out-of-sample predictions at $\mathrm{Da}_{s}=20{,}000$. The a priori-optimized neural network model is comparably accurate to the baseline $k$$\epsilon$ model in-sample but is unstable when evaluated for $\mathrm{Da}_{s}=20{,}000$. In both cases, the DPM-optimized ME0 model is more than one order of magnitude more accurate than each of the comparison models.
Figure 18: A posteriori RANS errors in predicted temperature (top) and cross-stream velocity (bottom) for $\mathrm{Da}_{s}=0$ jets using the $k$$\epsilon$ model, DPM-optimized ME0 model, direct DNS-evaluated $\mu_{t}^{k\epsilon}(Q_{\text{{dns}}};C_{\mu})$, and a priori-modeled $\mu_{t,ap}(Q;\theta)$.
$\mathrm{Da}_{s}=0$ $\mathrm{Da}_{s}=20{,}000$
Model MSE Rel. Error (%) MSE Rel. Error (%)
$k$$\epsilon$ 143.8 100.0 381.5 100.0
ME0 0.65 0.45 22.44 5.88
$\mu_{t}^{k\epsilon}(Q_{\text{{dns}}};C_{\mu})$ 39.81 27.7 145.43 38.1
$\mu_{t,ap}(Q;\theta)$ 150.85 104.9
Table 3: Cumulative MSE errors of RANS using the $k$$\epsilon$ model, the DPM-optimized ME0 model, direct DNS-evaluated $\mu_{t}^{k\epsilon}$, and the a priori-trained $\mu_{t,ap}$ model tested a posteriori for $\mathrm{Da}_{s}=0$ and $\mathrm{Da}_{s}=20{,}000$. The direct DNS-evaluated $\mu_{t}^{k\epsilon}$ is taken from $\mathrm{Da}_{s}=0$ for both testing Damköhler numbers. Relative errors are shown with respect to the baseline $k$$\epsilon$ model. The $\mathrm{Da}_{s}=0$-trained $\mu_{t,ap}(Q;\theta)$ model is unstable for $\mathrm{Da}_{s}=20{,}000$.
The superior performance of the DPM-optimized model over a priori methods can be attributed to the former’s inherent model consistency [32], which DPM achieves by accounting for feedback between the model and the governing PDEs during training. A priori optimization does not account for this feedback loop, which leads to poor predictive performance and potential instability in a posteriori applications [31, 33]. In the present study, the DPM approach is uniquely able to account for the influence of the RANS $k$ and $\epsilon$ variables on predictions, which is a significant advantage, as these variables are defined differently in RANS (in which they are the solution of PDEs) than they are in DNS (in which they are statistics of the instantaneous flow field). This is a feature of DPM that distinguishes it from the state of the art.
@@ 5.2 Comparison of DPM to a priori models
Another advantage of DPM is its ability to optimize over objective functions obtained from any PDE solution variables, including derived quantities such as spatially filtered and/or time-averaged statistics [33, 38]. This enables the DPM optimization procedure to target experimental data, from which the closure terms needed by a priori optimization may not be readily available, or even data from “higher-order” simulations such as direct simulation Monte Carlo solutions of the Boltzmann equation [36], which might lack obvious connections to the modeled closure terms. In contrast, a priori optimization must directly target the outputs of the closure model and hence is typically limited to DNS target data. Although DPM training is an order of magnitude more expensive than a priori methods, its advantages outweigh the training cost, while the inference cost remains comparable for both approaches.
@@
The main challenge of PDE-constrained optimization (not necessarily limited to DPM) is the cost of solving the optimization problem by iterating over the forward PDEs. DPM uses adjoint equations to obtain the gradients needed for optimization, which significantly alleviates the compute and memory requirements of non-graph-breaking alternatives (e.g., [51]). The adjoint equations can be complicated, especially for high-dimensional problems and detailed chemical kinetics. However, modern algorithmic differentiation frameworks, such as that provided by PyTorch [43], have greatly simplified the process of constructing the discrete-exact adjoint PDEs, which has made DPM accessible for large-scale flow simulations.